OpenStudy (mendicant_bias):

(Solving PDE's with Fourier Transforms) I'm wondering how exactly I should know where or when to input my initial conditions into a PDE that I've transformed, prompt below.

OpenStudy (mendicant_bias):

\[\gamma_{xx}+\gamma_{yy}=0, \ \ \ x>0, \ y>0.\] \[\gamma(x,y) \rightarrow 0 \ \text{as} \ \sqrt{x^2+y^2} \rightarrow \infty\]\[\gamma(0,y) =0, \ y > 0\]\[\gamma_y(x,0)=\frac{4x}{(4+x^2)^2}, \ x>0.\]

OpenStudy (mendicant_bias):

So what I'm doing is taking the Fourier Sine Treansform of both sides, but I still don't understand-how do I apply initial and boundary conditions to this problem in order to solve it?

OpenStudy (mendicant_bias):

@SithsAndGiggles @dan815

OpenStudy (mendicant_bias):

Or better said, how do I know *when* to apply boundary conditions? I can't just insert them wherever I feel like it.

OpenStudy (mendicant_bias):

And this problem has mixed conditions, right? Two dirichlet and one neumann?

OpenStudy (mendicant_bias):

(?)

OpenStudy (dan815):

hehe i dunno :) i gotta relearn PDE With FTs, I just crammed for my exam lol

OpenStudy (mendicant_bias):

Do you know who else I could ask about this stuff?

OpenStudy (dan815):

Kainui knows all this stuff

OpenStudy (mendicant_bias):

@Kainui

OpenStudy (mendicant_bias):

And thanks, I'm looking over problems 5 and 6 now.

OpenStudy (dan815):

he's sleeping you are better of sending him a PM the taggin stuff doesnt work well

OpenStudy (mendicant_bias):

@Kainui , when you have a PDE with mixed boundary conditions like this, should you apply an exponential transform? As far as I know (not sure) you (can? should?) only apply Fourier Sine Transforms when you have dirichlet conditions, and Fourier Cosine Transforms when you have neumann conditions.